OpenStudy (anonymous):

Solve 7x – 2y = –3 14x + y = 14

OpenStudy (anonymous):

x=5/7 y=4

OpenStudy (angela210793):

7x – 2y = –3 7x-2y=-3---> add both sides 35x=25-->x=25/35 14x + y = 14 /*(2)---? 28x+2y=28 x=5/7 7*(5/7)-2y=-3 5-2y=-3 -2y=-8 y=4

OpenStudy (anonymous):

Thanks so much guys :)

OpenStudy (angela210793):

yw :)